# The iterated Dirichlet process and applications to Bayesian inference

### v1

@@ 4.7 A proof without a simulation density
In order to prove that sequential imputation applies to the IDP, we must give a new proof based on a new assumption. We continue to let the simulation measure $\mathfrak{m}^{*}$ be defined as in Section 4.3. But we must drop Assumption 4.3, and consequently drop the assumption that $\mathfrak{m}^{*}$ has a density with respect to a product measure. We cannot drop densities altogether, though, since they are essential to defining the weight function.
@@
Under Assumption 4.6, we may let $f_{m}$ be a density of $(Y,{\bm{Z}}_{m})$ with respect to $\widetilde{\mathfrak{n}}^{M}\times\mathfrak{n}_{m}$. We adopt the same assumptions and notational conventions for $f_{m}$ as we did for $f$ in Section 4.4. For $(y,z)\in T\times S$, define
@@
[…]
@@
[…]
@@
almost surely. Note that $\gamma_{1}^{*}(Y,dz_{1})=\gamma_{1}(Y_{1},dz_{1})=f_{1}(z_{1}\mid Y_{1})\,\mathfrak{n}_{1}(dz_{1})$. Hence, starting with $k=M-1$ and iterating backwards to $k=1$, we obtain
@@
5 Sequential imputation for IDPs
In this section, we apply sequential imputation, in the form of Theorem 4.7, to an array of samples from an IDP. In Theorem 4.7, we take $Z=\bm{\mu}_{M}$ and we let $Y$ represent some observations we have made about the samples ${\bm{X}}_{MN}$.
In Section 5.1, we show that sequential imputation cannot, in general, be used on direct observations $Y={\bm{X}}_{MN}$. Starting in Section 5.2, therefore, we assume that $Y$ is a discrete function of ${\bm{X}}_{MN}$. Section 5.2 contains our main result, Theorem 5.2. This theorem shows how to use sequential imputation to compute $\mathcal{L}(\bm{\mu}_{M}\mid Y)$. The chief challenge is to construct the simulated row distributions, $\bm{\mu}_{M}^{*,k}$. According to (4.5), these should be constructed using the single-row conditional distributions, $\mathcal{L}({\mu_{m}\mid{\bm{Y}}_{m},\bm{\mu}_{m-1}})$. In Section 5.3, we compute $\mathcal{L}({\mu_{m}\mid{\bm{Y}}_{m},\bm{\mu}_{m-1}})$. In Section 5.4, we use these to generate $\bm{\mu}_{M}^{*,k}$. Then, in Section 5.5, we prove Theorem 5.2.
5.1 The impossibility of direct observations
A direct observation of the samples would be represented by taking $Y={\bm{X}}_{MN}$. In general, though, we cannot treat the case $Y={\bm{X}}_{MN}$ with sequential imputation. This is because Assumption 4.6 may fail. More specifically, if $Y={\bm{X}}_{MN}$ and the base measure $\varrho$ is not discrete, then Assumption 4.6 will fail, according to Proposition 5.1 below.
To prove Proposition 5.1, we begin by establishing two formulas that we will need later. Let $A\in\mathcal{S}^{n}$ and $B\in\mathcal{M}_{1}$. By Theorem 2.3,
$=E[1_{B}(\mu_{m})\mu_{m}^{n}(A)\mid\bm{\mu}_{m-1}].$
By (3.6), this gives
In particular, since $P(X_{mn}\in A)=E[\mu_{m}^{n}(A)]$, we have
Proposition 5.1.
Suppose $\varrho$ is not discrete. That is, there exists $B\in S$ such that $\varrho(B)>0$ and $\varrho(\{x\})=0$ for all $x\in B$. Then there do not exist $\sigma$-finite measures $\widetilde{\mathfrak{n}}_{1}$ and $\widetilde{\mathfrak{n}}_{2}$ on $S$ such that $\mathcal{L}(\xi_{11},\xi_{21})\ll\widetilde{\mathfrak{n}}_{1}\times\widetilde{\mathfrak{n}}_{2}$.
Proof.
Assume $\mathcal{L}(\xi_{11},\xi_{21})\ll\widetilde{\mathfrak{n}}_{1}\times\widetilde{\mathfrak{n}}_{2}$. Let $A=\{x\in S:\widetilde{\mathfrak{n}}_{1}(\{x\})>0\}$, so that $A$ is countable. Let $D=\{(x,x)\in S^{2}:x\in B\cap A^{c}\}$. Then
$(\widetilde{\mathfrak{n}}_{1}\times\widetilde{\mathfrak{n}}_{2})(D)=\int_{B\cap A^{c}}\widetilde{\mathfrak{n}}_{1}(\{x\})\,\widetilde{\mathfrak{n}}_{2}(dx)=0.$
Therefore, by hypothesis,
$P(\xi_{21}=\xi_{11},\xi_{11}\in B\cap A^{c})=P((\xi_{11},\xi_{21})\in D)=0.$ (5.3)
On the other hand, by Theorem 2.5 and (5.2),
$P(\xi_{21}\in C\mid\xi_{11})$ $=E[P(\xi_{21}\in C\mid\mu_{1},\xi_{11})\mid\xi_{11}]$
$=E[P(\xi_{21}\in C\mid\mu_{1})\mid\xi_{11}]$
$=E\bigg{[}\frac{1}{\kappa+1}(\kappa P(\xi_{21}\in C)+\mu_{1}(C))\;\bigg{|}\;\xi_{11}\bigg{]}$
$=\frac{\kappa}{\kappa+1}\,\varrho(C)+\frac{1}{\kappa+1}E[\mu_{1}(C)\mid\xi_{11}].$
By (3.5) and Theorem 3.1, this gives
$P(\xi_{21}\in C\mid\xi_{11})=\frac{\kappa}{\kappa+1}\,\varrho(C)+\frac{1}{\kappa+1}\left(\frac{\varepsilon}{\varepsilon+1}\,\varrho(C)+\frac{1}{\varepsilon+1}\,\delta_{\xi_{11}}(C)\right).$
Hence,
$\mathcal{L}(\xi_{21}\mid\xi_{11})=\frac{\tau}{\tau+1}\,\varrho+\frac{1}{\tau+1}\,\delta_{\xi_{11}},$
where $\tau=(\kappa+1)(\varepsilon+1)-1$. We therefore have
$P(\xi_{21}=\xi_{11},\xi_{11}\in B\cap A^{c})$ $=E[1_{B\cap A^{c}}(\xi_{11})P(\xi_{21}=\xi_{11}\mid\xi_{11})]$
$=E\bigg{[}1_{B\cap A^{c}}(\xi_{11})\bigg{(}\frac{\tau}{\tau+1}\,\varrho(\{\xi_{11}\})+\frac{1}{\tau+1}\bigg{)}\bigg{]}.$
Note that $\xi_{11}\in B$ implies $\varrho(\{\xi_{11}\})=0$. Thus,
$P(\xi_{21}=\xi_{11},\xi_{11}\in B\cap A^{c})=\frac{1}{\tau+1}P(\xi_{11}\in B\cap A^{c})=\frac{1}{\tau+1}\,\varrho(B\cap A^{c}).$
Since $A$ is countable,
$\varrho(B\cap A^{c})=\varrho(B)-\sum_{x\in A\cap B}\varrho(\{x\})=\varrho(B)>0,$
which contradicts (5.3). ∎
@@ 5.2 Sequential imputation with discrete observations
If $\varrho$ is discrete, then we may assume without loss of generality that $S$ is countable. Therefore, if $S$ is uncountable and $Y_{m}=X_{mN}=(\xi_{m1},\ldots,\xi_{mN})$, then Proposition 5.1 shows that Assumption 4.6 fails. In other words, if $S$ is uncountable and we wish to use sequential imputation, then we cannot observe the data $\xi_{ij}$ directly.
On the other hand, we can observe discrete functions of $\xi_{ij}$. This is because Assumption 4.6 is trivially satisfied whenever $Y$ is discrete. From an applied perspective, this is no restriction at all. Any real-world measurement will have limits to its precision, meaning that only a finite number of measurement outcomes are possible.
Theorem 5.2 below describes how to use sequential imputation to compute $\mathcal{L}(\bm{\mu}_{M}\mid Y)$ when our observations $Y$ are discrete. The proof of Theorem 5.2 will be given in Section 5.5. Before stating Theorem 5.2, we first establish some notation.
@@
Now fix $y=(y_{1},\ldots,y_{M})\in T^{M}$. Using $y$, we will construct a weighted simulation of $\bm{\mu}_{M}$, which is a pair $(t,\bm{u}_{M}^{*})$, where $t=\{t_{mi}:1\leq m\leq M,1\leq i\leq m\}$ is a triangular array of $[0,\infty)$-valued random variables and $\bm{u}_{M}^{*}=(u_{1}^{*},\ldots,u_{M}^{*})$ is a vector of $M_{1}$-valued random variables, all of which are independent of $Y$. The rows of $t$, which we denote by $t_{m}=(t_{m1},\ldots,t_{mm})$, are called the row weights of the weighted simulation, and the random measures $u_{m}^{*}$ are called the simulated row distributions. We construct $t_{m}$ and $u_{m}^{*}$ by recursion on $m$ as follows. Let
@@
The proof of Theorem 5.2 will be given in Section 5.5.
@@
Let $\Psi:M_{1}\to\mathbb{R}$ be continuous and bounded. By Theorem 2.5 and (3.6), we have
@@ 5.3 Conditioning on a single row
We will prove Theorem 5.2 by applying Theorem 4.7. To do this, we must, among other things, compute the conditional distribution $\gamma_{m}$ described in Section 4.3. This is done below, and the result is presented in (5.14).
@@
Let $\gamma$ be the probability kernel on the right-hand side of (5.14) and let $B\in\mathcal{M}_{1}$. We must show that $P(\mu_{m}\in B\mid{\bm{Y}}_{m},\bm{\mu}_{m-1})=\gamma({\bm{Y}}_{m},\bm{\mu}_{m-1},B)$. By Theorem 2.5, it suffices to show that $P(\mu_{m}\in B\mid Y_{m},\bm{\mu}_{m-1})=\gamma({\bm{Y}}_{m},\bm{\mu}_{m-1},B)$.
@@
Now let $C\in\sigma(Y_{m},\bm{\mu}_{m-1})$. Without loss of generality, we may assume that $C$ is of the form $C=\{Y_{m}\in D\}\cap\{\bm{\mu}_{m-1}\in F\}$ for some $D\subseteq T$ and $F\in\mathcal{M}_{1}^{m-1}$. Then (5.16) gives
@@ 5.4 Generating the simulations
To prove Proposition 5.5, we use the following explicit construction of $u_{m}^{*,k}(y)$. Define $H\subseteq\mathbb{R}^{m}$ by $H=[0,\infty)^{\infty}\setminus\{(0,\ldots,0)\}$. Let
@@
We first note that if $\mathcal{F}_{mk}=\sigma(U_{1k},\ldots,U_{mk},\lambda_{1k},\ldots,\lambda_{mk})$, then $t_{m}^{k}\in\mathcal{F}_{m-1,k}$ and $\theta(m,k)\in\mathcal{F}_{m-1,k}\vee\sigma(U_{mk})$. This follows from (5.18) and (5.19) by induction. Also, by (5.19), we have
@@
Hence, $u_{m}^{*,k}(y)\in\mathcal{F}_{mk}$. In particular, $U_{mk}$, $\lambda_{mk}$, and $\bm{u}_{m-1}^{*,k}(y)$ are independent.
Now let $B\in\mathcal{M}_{1}$ and $C\in\sigma(Y,\bm{\mu}_{m-1}^{*,k})$. Without loss of generality, we may assume that $C$ is of the form $C=\{Y\in D\}\cap\{\bm{\mu}_{m-1}^{*,k}\in F\}$ for some $D\subseteq T^{M}$ and $F\in\mathcal{M}_{1}^{m-1}$. We then have
@@ 5.5 Proof of the main result
For the other factors in (4.6), we use Theorem 2.5, (5.2), and (5.4) to obtain
@@ 6 Examples
In this section, we present four hypothetical applications to illustrate the use of the IDP. See https://github.com/jason-swanson/idp for the code used to generate the simulations.
The framework for each of these examples was described in Section 1. Namely, we consider a situation in which there are several agents, all from the same population. Each agent undertakes a sequence of actions. These actions are chosen according to the agent’s particular tendencies. Although different agents have different tendencies, there may be patterns in the population.
  • The future behavior of the agents we have observed.
  • The behavior of a new (unobserved) agent from the population.
We model this situation with an IDP. More specifically, let $S$ be a complete and separable metric space. Let $\xi$ be a row exchangeable array of $S$-valued random variables whose row distribution generator—the random measure $\varpi$ in Theorem 2.2—is an IDP. That is, $\varpi\sim\mathcal{D}(\kappa\mathcal{D}(\varepsilon\varrho))$, where $\kappa>0$ and $\varepsilon>0$ are the column and row concentration parameters, respectively, and $\varrho\in M_{1}$ is the base measure. The value of $\xi_{ij}$ represents the $j$th action of the $i$th agent. In this way, the space $S$ represents the set of possible actions that the agents may undertake.
@@
The outline of this section is as follows. In Section 6.1, we describe how Theorem 5.2 and Corollary 5.3 simplify in the case that $S$ is finite. After that, the remainder of the section is devoted to the examples. Our simplest example is in Section 6.2, and concerns the malfunctioning pressed penny machine described in Section 1. Section 6.3 presents a similar example, but with significantly more data. This is the same example treated in [12] (originally considered in [2]) and is concerned with the flicking of thumbtacks. Section 6.4 applies the IDP model to the analysis of Amazon reviews. The final example, found in Section 6.6, is about video game leaderboards. To prepare for that example, a custom prior distribution, which we call the “gamer” distribution, is presented in Section 6.5.
6.1 IDPs on a finite state space

### v2

@@ 3.6 A proof without a simulation density
We wish to apply sequential imputation to determine $\mathcal{L}(\bm{\mu}_{M}\mid{\bm{X}}_{MN})$, using the computable distributions (1.4). In this case, we would naturally take $Z_{m}=\mu_{m}$ and $Y_{m}=X_{mN}=(\xi_{m1},\xi_{m2},\ldots,\xi_{mN})$. In [11], Liu alleged to do exactly this in the special case $S=\{0,1\}$, using the results in [10] as his justification.
Unfortunately, sequential imputation—as it is presented in Theorem 3.4—does not apply in this case. Namely, Assumption 3.3 is not satisfied. In fact, it is straightforward to verify that, as long as $\varrho$ is not a point mass, the vector $Z=\bm{\mu}_{m}$ has no joint density with respect to any product measure.
Hence, the proof of Theorem 3.4, which is a rigorous presentation of the proof in [10], does not justify the use of sequential imputation in this setting. This includes not only the general setting that we are working with, but also the special case $S=\{0,1\}$ that was treated in [11].
In this section, we give a new proof of (3.6), in which we do not require the joint densities of Assumption 3.3. In doing so, we retroactively justify the results in [11], and also lay the foundations for applying sequential imputation to the NDP on an arbitrary state space $S$.
We continue to let the simulation measure $\mathfrak{m}^{*}$ be defined as in Section 3.3, but we must drop the assumption that $\mathfrak{m}^{*}$ has a density with respect to a product measure. We cannot drop densities altogether, though, since they are essential to defining the weight function.
@@
Under Assumption 3.5, we may let $f_{m}$ be a density of $(Y,{\bm{Z}}_{m})$ with respect to $\widetilde{\mathfrak{n}}^{M}\times\mathfrak{n}_{m}$. We adopt the same assumptions and notational conventions for $f_{m}$ as we did for $f$ in Section 3.4. For $(y,z)\in T^{M}\times S^{M}$, define
@@
$\mathcal{L}(\widetilde{Z}^{K}\mid Z^{*,1},\ldots,Z^{*,K},Y)\to\mathcal{L}(Z\mid Y)\quad\text{a.s.}$
@@
$w(Y,\cdot)=f_{M}(Y)\,\frac{d\mathcal{L}(Z\mid Y)}{d\mathfrak{m}^{*}(Y)}\quad\text{a.s.}$ (3.8)
@@
almost surely. Note that $\gamma_{1}^{*}(Y,dz_{1})=\gamma_{1}(Y_{1},dz_{1})=f_{1}(z_{1}\mid Y_{1})\,\mathfrak{n}_{1}(dz_{1})$. Hence, starting with k=M1k=M-1 and iterating backwards to $k=1$, we obtain
@@
4 Sequential imputation for the NDP
In this section, we apply sequential imputation, in the form of Theorem 3.6, to an array of samples from an NDP. In Theorem 3.6, we take $Z_{m}=\mu_{m}$ and we let $Y_{m}$ represent some observations we have made about the samples $X_{mN}$.
A direct observation of the samples would be represented by taking $Y_{m}=X_{mN}$. In general, though, we cannot treat the case $Y_{m}=X_{mN}$ with sequential imputation. This is because Assumption 3.5 may fail. Part of Assumption 3.5 is that $Y=(Y_{1},\ldots,Y_{M})$ has a joint density with respect to some product measure. But when $Y_{m}=X_{mN}$, this fails even in the case $M=2$ and $N=1$. More specifically, it is straightforward to verify that if $\varrho$ is not discrete, then $(\xi_{11},\xi_{21})$ has no joint density with respect to any product measure.
On the other hand, we can observe discrete functions of $X_{mN}$. This is because Assumption 3.5 is trivially satisfied whenever $Y$ is discrete. From an applied perspective, this is no restriction at all. Any real-world measurement will have limits to its precision, meaning that only a finite number of measurement outcomes are possible. We therefore assume, from this point forward, that $Y_{m}$ is a discrete function of $X_{mN}$.
Section 4.1 contains our main result, Theorem 4.1. This theorem shows how to use sequential imputation to compute $\mathcal{L}(\bm{\mu}_{M}\mid Y)$. The chief challenge is to construct the simulated row distributions, $\bm{u}_{M}^{*,k}$. According to (3.5), these should be constructed using the single-row conditional distributions, $\mathcal{L}({\mu_{m}\mid{\bm{Y}}_{m},\bm{\mu}_{m-1}})$. In Section 4.2, we compute $\mathcal{L}({\mu_{m}\mid{\bm{Y}}_{m},\bm{\mu}_{m-1}})$. In Section 4.3, we use these to generate $\bm{u}_{M}^{*,k}$. Finally, in Section 4.4, we give the proof of Theorem 4.1.
@@ 4.1 Sequential imputation with discrete observations
Now fix $y=(y_{1},\ldots,y_{M})\in T^{M}$. Using $y$, we will construct a weighted simulation of $\bm{\mu}_{M}$, which is a pair $(t,\bm{u}_{M}^{*})$, where $t=\{t_{mi}:1\leq m\leq M,1\leq i\leq m\}$ is a triangular array of $0,\infty)-$valued random variables and $\bm{u}_{M}^{*}=(u_{1}^{*},\ldots,u_{M}^{*})$ is a vector of $M_{1}$-valued random variables, all of which are independent of $Y$. The rows of $t$, which we denote by $t_{m}=(t_{m1},\ldots,t_{mm})$, are called the row weights of the weighted simulation, and the random measures $u_{m}^{*}$ are called the simulated row distributions. We construct $t_{m}$ and $u_{m}^{*}$ by recursion on $m$ as follows. Let
@@
The above is what the distribution of $\mu_{m}$ would be if we had only observed the row $y_{m}$. (This is a consequence of (2.8).)
@@
The proof of Theorem 4.1 will be given in Section $4.$
@@
Let $\Psi:M_{1}\to\mathbb{R}$ be continuous and bounded. Using (2.3) and the fact that $\mu_{M+1}$ and $Y$ are conditionally independent given $\bm{\mu}_{M}$, we have
@@ 4.2 Conditioning on a single row
We will prove Theorem 4.1 by applying Theorem 3.6. To do this, we must, among other things, compute the conditional distribution $\gamma_{m}$ described in Section 3.3. This is done below, and the result is presented in (4.14). In order to derive this result, we begin by establishing some formulas that we will need later.
Note that $\sigma(\prod_{i=1}^{m}\mu_{i}^{n})\subseteq\sigma(\varpi,\bm{\mu}_{m})\subseteq\sigma(\mu)$. Also, since $\{\xi_{ij}\}$ is row exchangeable, we have $\mathcal{L}({\bm{X}}_{mn}\mid\mu)=\prod_{i=1}^{m}\mu_{i}^{n}$. Therefore,
$\mathcal{L}({\bm{X}}_{mn}\mid\varpi,\bm{\mu}_{m})=\mathcal{L}({\bm{X}}_{mn}\mid\bm{\mu}_{m})=\prod_{i=1}^{m}\mu_{i}^{n}.$ (4.10)
Now let $A\subseteq S^{n}$ and $B\subseteq M_{1}$ be Borel. By (4.10),
$=E[1_{B}(\mu_{m})\mu_{m}^{n}(A)\mid\bm{\mu}_{m-1}].$
By (2.3), this gives
In particular, since $P(X_{mn}\in A)=E[\mu_{m}^{n}(A)]$, we have
@@
Let $\gamma$ be the probability kernel on the right-hand side of (4.14) and let $B\subseteq M_{1}$ be Borel. We must show that $P(\mu_{m}\in B\mid{\bm{Y}}_{m},\bm{\mu}_{m-1})=\gamma({\bm{Y}}_{m},\bm{\mu}_{m-1},B)$. Since ${\bm{X}}_{m-1,N}$ and $\mu_{m}$ are conditionally independent given $\bm{\mu}_{m-1}$, it suffices to show that $P(\mu_{m}\in B\mid Y_{m},\bm{\mu}_{m-1})=\gamma({\bm{Y}}_{m},\bm{\mu}_{m-1},B)$.
@@
Now let $C\in\sigma(Y_{m},\bm{\mu}_{m-1})$. Without loss of generality, we may assume that $C$ is of the form $C=\{Y_{m}\in D\}\cap\{\bm{\mu}_{m-1}\in F\}$ for some $D\subseteq T$ and some Borel $F\subseteq M_{1}^{m-1}$. Then (4.16) gives
@@ 4.3 Generating the simulations
To prove Proposition 4.4, we use the following explicit construction of $u_{m}^{*,k}(y)$. Define $H\subseteq\mathbb{R}^{m}$ by $H=[0,\infty)^{m}\setminus\{(0,\ldots,0)\}$. Let
@@
We first note that if $\mathcal{F}_{mk}=\sigma(U_{1k},\ldots,U_{mk},\lambda_{1k},\ldots,\lambda_{mk})$, then $t_{m}^{k}$ is $\mathcal{F}_{m-1,k}$-measurable and $\theta(m,k)$ is $\mathcal{F}_{m-1,k}\vee\sigma(U_{mk})$-measurable. This follows from (4.18) and (4.19) by induction. Also, by (4.19), we have
@@
Hence, $u_{m}^{*,k}(y)$ is $\mathcal{F}_{mk}$-measurable. In particular, $U_{mk}$, $\lambda_{mk}$, and $\bm{u}_{m-1}^{*,k}(y)$ are independent.
Now let $B\subseteq M_{1}$ be Borel and let $C\in\sigma(Y,\bm{\mu}_{m-1}^{*,k})$. Without loss of generality, we may assume that $C$ is of the form $C=\{Y\in D\}\cap\{\bm{\mu}_{m-1}^{*,k}\in F\}$ for some $D\subseteq T^{M}$ and some Borel $F\subseteq M_{1}^{m-1}$. We then have
@@ 4.4 Proof of the main result
For the other factors in (3.7), we use (4.12) and (4.1), and the fact that ${\bm{Y}}_{m}$ and $Y_{m+1}$ are conditionally independent given $\bm{\mu}_{m}$ to obtain
@@ 5 Examples
In this section, we present four hypothetical applications to illustrate the use of the Theorem 4.1 and Corollary 4.2. See https://github.com/jason-swanson/ndp for the code used to generate the simulations.
The framework for each of these examples was described in Section 1. In that framework, we interpret $\xi_{ij}$ as the $j$th action of the $i$th agent. The space $S$ is therefore the set of possible actions.
@@
The outline of this section is as follows. In Section 5.1, we describe how Theorem 4.1 and Corollary 4.2 simplify in the case that $S$ is finite. After that, the remainder of the section is devoted to the examples. Our simplest example is in Section 5.2, and concerns a malfunctioning pressed penny machine. Section 5.3 presents a similar example, but with significantly more data. This is the same example treated in [11] (originally considered in [3]) and is concerned with the flicking of thumbtacks. Section 5.4 applies the NDP model to the analysis of Amazon reviews. The final example, found in Section 5.6, is about video game leaderboards. To prepare for that example, a custom prior distribution, which we call the “gamer” distribution, is presented in Section 5.5.
5.1 The case of a finite state space